# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ III LLM-Driven AI Agents
In this section, we review the entire lifetime from LLM to LLM-driven AI agent. Our goal is to help beginners to quickly figure out agents, their characteristics, relationships, and applications.
@@ III-A Large Language Model
Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [200]. Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [311], logic reasoning [212, 339, 281], code generation [325], and translation [209]. These remarkable performances can be attributed to two major factors. One is that LLMs are built upon a powerful architecture known as the Transformer [258], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input. The other key factor, perhaps the most important one, is the massive scales of LLMs that far exceeds traditional AI models. When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [280], referring to unexpected capabilities that do not appear in smaller models. As shown in Table II, the parameter scale of a LLM can be hundreds or thousands of times that of traditional AI models.
@@ III-B LLM-Driven AI Agents
Figure 3 illustrates a typical architecture of LLM-driven agents. Different from LLMs that mainly act as chat bots and do not possess professional ability in specific domains, agents are designed to automatically help human to finish specialized tasks. To this end, agents are equipped with multiple modules to become more all-powerful. As shown in Figure 3, there are usually five modules in agents: perception, memory, tools, reasoning, and action.
  • Perception module. To automatically finish a specified task, agents need the ability to perceive the real-world environment. For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [183, 179]. The type of perception ability depends on the domain for which the agent is designed. For instance, an autonomous driving agent need the ability of visual or radar perception [244, 302], while a code-generating agent may not require such functions [106, 114].
  • Memory module. The processing of real-world tasks also require the strong ability of memory. Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [92, 180, 309]. This usually require external storage resources to assist the brain, such as database or memory sharing [75, 77]. In contrast, LLMs do not have such an excellent memory ability. Their memory is short-term, which only lasts for rounds of conversations [266, 340].
  • Reasoning and planing module. LLM acts as the brain of agents due to its excellent ability of reasoning and planing. It intercepts the instructions from users and automatically decompose the received task into multiple feasible steps [273, 117, 129, 240]. Then, it selects a best plan from different candidates [126, 336, 103]. Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [344, 257, 221, 286]. For example, when the autonomous driving module finds that the barrier is closer, it will changes the plan to slow down or detour.
@@
  • Action module. The action module is the core hub for the interaction with the environment. It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [272, 326]. This module ensures the executability of instructions through structured output control. For example, it immediately stops generating when LLM generate a complete action description to avoid redundant output interfering with subsequent parsing.
By integrating the above modules, agents establishes a closed-loop system that achieves a full chain of perception-decision-action-feedback. As a result, agents achieve unprecedented ability in automatically finishing domain-specific tasks, being closer to the ultimate form of AI that human expects.
@@ III-C Comparison Between Agents and LLMs
Table III illustrates the advantages of agents over LLMs on different metrics. Overall, agents have many advantages over LLMs except security.
  • High Autonomy. LLMs can only passively react to the user prompts and then generate responses. They are unable to plan or execute tasks independently. Besides, the response quality highly rely on the prompt skill [283, 80, 191, 167, 63, 29, 343], which seriously affects the user experience. In contrast, agents possess independent capabilities for task decomposition, strategy adjustment, and external tool invocation, which breaks through the passive mode of LLMS and is highly autonomous.
  • Flexible Multimodal interaction. LLMs have limited capability of handling multimodal inputs, such as text and pictures [227, 319, 323, 343, 144, 346]. Besides, their outputs are also mainly single-modal (e.g., text-only or picture-only), lacking the ability to actively invoke tools to perform physical actions or generate multimodal content. In contrast, agents overcome this drawbacks by deploying multimodal perception frameworks and tool invocation interfaces. They can realize interactions with complex environments, including vision, text, voice, and other physical elements.
  • Abundant Tool invocation. LLMS usually passively invoke a single tool (such as Function Calling [202]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [318]). In contrast, agents have the ability of active decision-making. They can independently select, combine and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [97].
@@
  • Dynamic adaptability. Essentially, LLMs are static models whose knowledge is fixed at the training phase. Although techniques such as fine-tuning [315, 164, 105] or model editing [303, 270, 322, 265, 162] reduce the training cost significantly, LLMs still cannot adapt to real-time events well. In contrast, agents are equipped with techniques like on-line web search, database query, or real-time sensors, which make them able to dynamically adapt to the changes of real-time environments and information.
  • Stronger Collaboration ability. LLMs lack enough collaboration ability when handling complex tasks. First, LLMs cannot interact with tools well, they can only access limited external assistance via simple API. Second, different LLMs lack effective cooperation mechanisms. In contrast, agents have designs for multi-agent collaboration. For example, MCP enables agents to use unified integration of external tools, and A2A allows agents from different enterprises to cooperatively finish a task.
  • Worse Security. Agents have WORSE security than LLMs, which is a major weakness of agents. This is because LLMs are only capable of outputting text. Even if the outputs contain illegal or discriminatory contents, their influences to the real world are limited. In contrast, since agents are endowed with the ability to invoke tools, they can cause substantial damages to the real world, including but not limited to maliciously/wrongly operating machines, poisoning databases, and paralyzing the system. As a result, it is necessary to concentrating more on the security of agents.
@@ III-D Agent Applications
Due to the strong advantages that agents have shown, related applications are booming. They span multiple domains, from scientific research to engineering systems and social services. Since the application of agents is not the focus of this paper, we will present an brief overview of their practical use cases to illustrate the rapid popularization of agent.
@@
Technical and Engineering Systems. Agents play a growing role in engineering domains, improving automation, system and software intelligence. For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [262, 33, 115, 104, 128, 176]. Besides, agents are also popular in game development and simulation [189, 241]. Embodied intelligence is also another hot topic [32, 182].
Social Governance and Public Services Agent are increasingly deployed in sectors focused on public service and human welfare. For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [254, 111, 187, 253]. Besides, other fields, such as financial services [155, 300, 201, 310, 89, 88, 67], education [269, 245, 181, 55, 59, 194], and healthcare [31, 279, 68, 225, 19, 113, 271, 148, 327], are also actively integrating agents into their respective practices.
@@ III-E Takeaways
Agents show multiple advantages over LLMs on multiple metrics, such as richer perception ability, stronger learning ability, and higher adaptivity. Now, agents are quickly deployed in various domains, providing unprecedented assistance to different groups. Especially, to improve the service quality, agents are evolving towards refinement to obtain professional skill in a small domain, no longer pursuing the comprehensive capabilities like LLM. In contrast, LLMs are more like an intermediate transitional form of the future intelligence, while agents are the next stage development direction of artificial intelligence. It can be foreseen that they will ultimately become indispensable components of future production ecosystems and daily life. However, agents show worse security than LLMs due to its capability of executing tools. As a result, studying the security of agent communication is significant to the AI ecosystem.
@@ IV-A Motivation: The Demand for Agent Communication
Although the advantages of agents in various fields have become increasingly obvious, their development has also encountered new obstacles, which has given rise to the demand for agent communication.
Conflicting development trends. The first reason derives from the fundamental conflict between the refined development of agents and the abstract demands of users. With the in-depth and specialized evolution of agents towards vertical fields (such as medical diagnostic agents, financial risk control agents, and industrial control agents), their capability boundaries are becoming increasingly refined. However, users’ usage habits exhibit opposite characteristics: they tend to input simple and abstract instructions (such as “plan a cross-border travel”) to trigger the execution of complex tasks. It is hard for a domain-specific agent to finish such abstract instructions independently. Besides, this trend of users’ habits is hard to be reversed. This is because people always prefer applications that are easy to operate rather than those that require cumbersome usage steps. The latter is at a disadvantage in the market competition. Usually, for each additional operation step, the user churn rate increases by 10% - 20%. Therefore, agents should not only NOT ask users to make changes, but also cater to such demands of users, which is contrary to to the development direction of agents.
Closed ecosystem. The current mainstream multi-agent systems adopt a closed ecosystem design and rely on private interaction mechanisms, forming a rigid technical barrier. This development paradigm severely limits the dynamic collaboration capabilities with external systems. For example, external agents cannot be actively discovered or invoked, making it extremely hard for cross-platform collaboration. Besides, such closed ecosystem further triggers a decline in systemic effectiveness. On one hand, agents loses scalability due to binding private tools. On the other hand, the obstruction of cross-platform knowledge sharing suppresses the intelligence of agents. As a result, it is necessary to propose agent communication mechanisms that integrate agents from different communities.
@@ IV-B Agent Communication Definition
When an agent complete tasks, it conducts multimodal information exchange and dynamic behavior coordination with diversified elements through standardized protocol frameworks, and finally return the results to the user. The communication behaviors in this process all belong to agent communication.
@@
  • Agent communication is task-driven. All types of agent communication must be invoked under the condition that users assigned a task. Although in some scenarios, the instructions received by agents are from another agent instead of users, these invoking processes can also be traced back to an original user instruction finally. Therefore, such communication is also regarded as agent communication. In contrast, for example, when no user tasks are generated, the update of the database or the synchronization of the distributed databases is not regarded as agent communication.
  • One of the communication objects must be an agent. Agent can communicate with different elements, such as tools, users, or other agents. As long as one of the communication objects is agent, this communication is regarded as agent communication. In contrast, for example, if users directly query the database to refine their instructions before submitting to agents, this user-database interaction is not regarded as agent communication. If the invoked tool call other tools (e.g., a computation tools calls other libraries), this process is not agent communication.
@@ IV-C1 User-Agent Interaction
User-agent Interaction refers to the interaction process in which agents receive user instructions and feed back execution results to the user. As shown in Figure 4, the user issues a task to an agent in step 1, i.e., make a travel plane to Beijing. The agent conducts a series of actions to complete this task and finally sends the result to the user in step 7. Please note that the interaction process between users and agents is fundamentally similar to interacting with LLMs. Therefore, we adopt the term interaction rather than communication.
@@ IV-C2 Agent-Agent Communication
Agent-agent communication is the communication process in which two or more agents conduct negotiation, task decomposition, sub-task allocation, and result aggregation for the collaborative completion of user-assigned tasks through standardized collaboration protocols. In Figure 4, the agent decomposes the travel task and assign sub-tasks (step 3). For example, this task is decomposed into searching scenic spots, checking weather, booking ticket, and hotel reservation, and each sub-task is conducted by an independent agent. Then, the agent seeks proper agents on the Internet and assigning these tasks to them (step 4). These agents will finish the received tasks and return the results to the original agent (step 6).
@@ IV-C3 Agent-Environment Communication
Agent-environment communication refers to the communication process in which agents conduct interactions with environmental entities (i.e., tools, equipments, and any other external elements helpful for task execution) through standardized protocols to complete user tasks. In Figure 4, before assigning tasks to other agents, the original agent queries the weather of Beijing through on-line search (step 2), which is a typical agent-environment communication case. Besides, other agents can also complete sub-tasks with the help of environmental tools. For example, in step 5, the travel agent searches the popular tourist spots through its database or searching on-line blogs.
@@
IV-D Three-layered Organization of Agent Communication Discussion
As shown in Figure 5, in the following paper, we will use a three-layered architecture to discuss agent communication and related security. In Section V, we will introduce user-agent interaction, Section VI is responsible for showing agent-agent communication, and Section VII is the content related to agent-environment communication. For each section, we still use a three-layered organization: firstly, we show related communication protocols; secondly, we analyze related security risks; finally, we outlook on defense countermeasures. Please note that the risks/defenses we focus on are specific for agent communication instead of LLMs. This is because agent and agent communication still have significant differences from LLMs (as we have discussed in Section III). For example, in the agent collaboration process, the accumulation of tiny benign deviation on each step may lead to an intolerable risk in the final result. Miehling et al. [192] also point out that over-focusing on the capabilities of a single model can result in neglecting the emergent behaviors at the system level and underestimating the true risks. As a result, for LLM-targeted attacks that have valuable inspiration to agent security, we will discuss the typical representatives of them instead of listing all papers out exhaustively.

### v2

@@ III LLM-Driven AI Agents
In this section, we review the entire lifetime from LLM to LLM-driven AI agent. Our goal is to help beginners quickly figure out agents, their characteristics, relationships, and applications.
@@ III-A Large Language Model
Large Language Model (LLM) is a new type of artificial intelligence (AI) model trained on large-scale text corpora to understand and generate human language [211]. Once it came out, LLMs have demonstrated unprecedented capabilities across a wide range of domains, including but not limited to natural language understanding and generating [330], logic reasoning [225, 356, 300], code generation [344], and translation [222]. These remarkable performances can be attributed to two major factors. One is that LLMs are built upon a powerful architecture known as the Transformer [276], which effectively models and captures contextual dependencies between tokens and dynamically weighs the importance of different parts of the input. The other key factor, perhaps the most important one, is the massive scales of LLMs that far exceed traditional AI models. When model parameters surpass certain thresholds, LLMs exhibit emergent abilities [299], referring to unexpected capabilities that do not appear in smaller models. As shown in Table II, the parameter scale of an LLM can be hundreds or thousands of times that of traditional AI models.
@@ III-B LLM-Driven AI Agents
Figure 3 illustrates a typical architecture of LLM-driven agents. Different from LLMs that mainly act as chat bots and do not possess professional ability in specific domains, agents are designed to automatically help humans to finish specialized tasks. To this end, agents are equipped with multiple modules to become more all-powerful. As shown in Figure 3, there are usually five modules in agents: perception, memory, tools, reasoning, and action.
  • Perception module. To automatically finish a specified task, agents need the ability to perceive the real-world environment. For example, the autonomous driving agent needs to sense road conditions in real time so as to take actions such as avoiding, driving, or braking [195, 190]. The type of perception ability depends on the domain for which the agent is designed. For instance, an autonomous driving agent needs the ability of visual or radar perception [258, 321], while a code-generating agent may not require such functions [116, 124].
  • Memory module. The processing of real-world tasks also requires a strong memory. Agents need to have long-term memory to store complex instructions, knowledge, environment interaction history, or other data that may be required in future steps [102, 191, 328]. This usually requires external storage resources to assist the brain, such as databases or memory sharing [80, 82]. In contrast, LLMs do not have such excellent memory ability. Their memory is short-term, which only lasts for rounds of conversations [284, 357].
  • Reasoning and planning module. LLM acts as the brain of agents due to its excellent capability of reasoning and planning. It intercepts the instructions from users and automatically decomposes the received task into multiple feasible steps [291, 127, 139, 253]. Then, it selects the best plan from different candidates [136, 354, 113]. Besides, it also revises strategies based on environmental feedback, mitigating errors like code bugs or logical inconsistencies [361, 275, 234, 305]. For example, when the autonomous driving module finds that the barrier is closer, it will change the plan to slow down or detour.
@@
  • Action module. The action module is the core hub for interaction with the environment. It is responsible for converting the decisions made by LLMs into executable physical or digital operations and obtaining feedback [290, 345]. This module ensures the executability of instructions through structured output control. For example, it immediately stops generating when LLMs generate a complete action description to avoid redundant output interfering with subsequent parsing.
By integrating the above modules, agents establish a closed-loop system that achieves a full chain of perception-decision-action-feedback. As a result, agents achieve unprecedented ability in automatically finishing domain-specific tasks, being closer to the ultimate form of AI that human expects.
@@ III-C Comparison Between Agents and LLMs
Table III illustrates the advantages of agents over LLMs on different metrics. Overall, agents have many advantages over LLMs, except for security.
  • High Autonomy. LLMs can only passively react to the user prompts and then generate responses. They are unable to plan or execute tasks independently. Besides, the response quality highly relies on the prompt skill [302, 85, 203, 178, 68, 33, 360], which seriously affects the user experience. In contrast, agents possess independent capabilities for task decomposition, strategy adjustment, and external tool invocation, which breaks through the passive mode of LLMS and is highly autonomous.
  • Flexible Multimodal interaction. LLMs have limited capability of handling multimodal inputs, such as text and pictures [240, 338, 342, 360, 155, 363]. Besides, their outputs are also mainly single-modal (e.g., text-only or picture-only), lacking the ability to actively invoke tools to perform physical actions or generate multimodal content. In contrast, agents overcome these drawbacks by deploying multimodal perception frameworks and tool invocation interfaces. They can realize interactions with complex environments, including vision, text, voice, and other physical elements.
  • Abundant Tool invocation. LLMS usually passively invoke a single tool (such as Function Calling [213]) through predefined API interfaces and can only perform fixed operations as instructed (e.g., calling the weather API to answer queries [337]). In contrast, agents have the capability of active decision-making. They can independently select, combine, and dynamically adjust multiple tools, such as connecting crawlers, databases, and visualization tools, to generate responses [107].
@@
  • Dynamic adaptability. Essentially, LLMs are static models whose knowledge is fixed at the training phase. Although techniques such as fine-tuning [334, 175, 115] or model editing [322, 288, 341, 283, 171] reduce the training cost significantly, LLMs still cannot adapt to real-time events well. In contrast, agents are equipped with techniques like online web search, database query, or real-time sensors, which enable them to dynamically adapt to the changes of real-time environments and information.
  • Stronger Collaboration ability. LLMs lack enough collaboration ability when handling complex tasks. First, LLMs cannot interact with tools well; they can only access limited external assistance via simple APIs. Second, different LLMs lack effective cooperation mechanisms. In contrast, agents have designs for multi-agent collaboration. For example, MCP enables agents to use unified integration of external tools, and A2A allows agents from different enterprises to cooperatively finish a task.
  • Worse Security. Agents have WORSE security than LLMs, which is a major weakness of agents. This is because LLMs are only capable of outputting text. Even if the outputs contain illegal or discriminatory content, their influence on the real world is limited. In contrast, since agents are endowed with the ability to invoke tools, they can cause substantial damages to the real world, including but not limited to maliciously/wrongly operating machines, poisoning databases, and paralyzing the system. As a result, it is necessary to concentrate more on the security of agents.
@@ III-D Agent Applications
Due to the strong advantages that agents have shown, related applications are booming. They span multiple domains, from scientific research to engineering systems and social services. Since the application of agents is not the focus of this paper, we will present a brief overview of their practical use cases to illustrate the rapid popularization of agents.
@@
Technical and Engineering Systems. Agents play a growing role in engineering domains, improving automation, systems, and software intelligence. For example, agents are widely used in software engineering, assisting in code generation, bug localization, verification, and system configuration [280, 37, 125, 114, 138, 187]. Besides, agents are also popular in game development and simulation [201, 255]. Embodied intelligence is also another hot topic [36, 194].
Social Governance and Public Services Agents are increasingly deployed in sectors focused on public service and human welfare. For example, agents are now widely used in the legal field to help draft contracts, review legal documents, check compliance rules, and analyze cases [270, 121, 199, 269]. Besides, other fields, such as financial services [164, 319, 212, 329, 98, 97, 73], education [287, 259, 193, 59, 64, 206], and healthcare [35, 298, 74, 238, 23, 123, 289, 159, 346], are also actively integrating agents into their respective practices.
@@ III-E Takeaways
Agents show multiple advantages over LLMs on multiple metrics, such as richer perception ability, stronger learning ability, and higher adaptivity. Now, to improve the service quality, agents are evolving towards refinement to obtain professional skill in a small domain, no longer pursuing the comprehensive capabilities like LLM. LLMs are more like an intermediate transitional form of the future intelligence, while agents are the next stage of development direction of artificial intelligence. It can be foreseen that they will ultimately become indispensable components of future production ecosystems and daily life. However, agents show worse security than LLMs due to their capability of executing tools. As a result, studying the security of agent communication is significant to the AI ecosystem.
@@ IV-A Motivation: The Demand for Agent Communication
Although the advantages of agents in various fields have become increasingly obvious, their development has also encountered new obstacles, which have given rise to the demand for agent communication.
Conflicting development trends. The first reason derives from the fundamental conflict between the refined development of agents and the abstract demands of users. With the in-depth and specialized evolution of agents towards vertical fields (such as medical diagnostic agents, financial risk control agents, and industrial control agents), their capability boundaries are becoming increasingly refined. However, users’ usage habits exhibit opposite characteristics: they tend to input simple and abstract instructions (such as “plan a cross-border travel”) to trigger the execution of complex tasks. It is hard for a domain-specific agent to finish such abstract instructions independently. Besides, this trend of users’ habits is hard to reverse. This is because people always prefer applications that are easy to operate rather than those that require cumbersome usage steps. The latter is at a disadvantage in the market competition. Usually, for each additional operation step, the user churn rate increases by 10% - 20%. Therefore, agents should not only NOT ask users to make changes, but also cater to such demands of users, which is contrary to the development direction of agents.
Closed ecosystem. The current mainstream multi-agent systems adopt a closed ecosystem design and rely on private interaction mechanisms, forming a rigid technical barrier. This development paradigm severely limits the dynamic collaboration capabilities with external systems. For example, external agents cannot be actively discovered or invoked, making it extremely hard for cross-platform collaboration. Besides, such a closed ecosystem further triggers a decline in systemic effectiveness. On one hand, agents lose scalability due to binding private tools. On the other hand, the obstruction of cross-platform knowledge sharing suppresses the intelligence of agents. As a result, it is necessary to propose agent communication mechanisms that integrate agents from different communities.
@@ IV-B Agent Communication Definition
When an agent completes tasks, it conducts multimodal information exchange and dynamic behavior coordination with diversified elements through standardized protocol frameworks, and finally returns the results to the user. The communication behaviors in this process all belong to agent communication.
@@
  • Agent communication is task-driven. All types of agent communication must be invoked under the condition that users assign a task. Although in some scenarios, the instructions received by agents are from another agent instead of users, these invoking processes can also be traced back to an original user instruction. Therefore, such communication is also regarded as agent communication. In contrast, for example, when no user tasks are generated, the update of the database or the synchronization of the distributed databases is not regarded as agent communication.
  • One of the communication objects must be an agent. Agents can communicate with different elements, such as tools, users, or other agents. As long as one of the communication objects is an agent, this communication is regarded as agent communication. In contrast, for example, if users directly query the database to refine their instructions before submitting to agents, this user-database interaction is not regarded as agent communication. If the invoked tool calls other tools (e.g., a computation tool calls other libraries), this process is not agent communication.
@@ IV-C1 User-Agent Interaction
User-agent Interaction refers to the interaction process in which agents receive user instructions and feed back execution results to the user. As shown in Figure 4, the user issues a task to an agent in step 1, i.e., make a travel plan to Beijing. The agent conducts a series of actions to complete this task and finally sends the result to the user in step 7. Please note that the interaction process between users and agents is fundamentally similar to interacting with LLMs. Therefore, we adopt the term interaction rather than communication.
Figure 5: Taxonomy of our survey of agent communication protocols, security risks, and defense countermeasures. The dotted line indicates that a defense can mitigate the risk pointed to by the arrow.
@@ IV-C2 Agent-Agent Communication
Agent-agent communication is the communication process in which two or more agents conduct negotiation, task decomposition, sub-task allocation, and result aggregation for the collaborative completion of user-assigned tasks through standardized collaboration protocols. In Figure 4, the agent decomposes the travel task and assigns sub-tasks (step 3). For example, this task is decomposed into searching scenic spots, checking the weather, booking a ticket, and hotel reservation, and each sub-task is conducted by an independent agent. Then, the agent seeks proper agents on the Internet and assigns these tasks to them (step 4). These agents will finish the received tasks and return the results to the original agent (step 6).
@@ IV-C3 Agent-Environment Communication
Agent-environment communication refers to the communication process in which agents conduct interactions with environmental entities (e.g., tools, knowledge databases, and any other external resources helpful for task execution) through standardized protocols to complete user tasks. In Figure 4, before assigning tasks to other agents, the original agent queries the weather of Beijing through online search (step 2), which is a typical agent-environment communication case. Besides, other agents can also complete sub-tasks with the help of environmental tools. For example, in step 5, the travel agent searches the popular tourist spots through its database or searches online blogs.
@@
IV-D Organization of the Following Sections
As shown in Figure 5, in the following sections, we will discuss agent communication and related security.
  • In Section V, we will introduce user-agent interaction. Specifically, the risks in this process are divided into two categories: risks from malicious users against benign agents (Section V-B) and risks from compromised agents against benign users (Section V-C). The possible defense countermeasures against malicious users are discussed in Section V-D, while those against malicious agents are discussed uniformly in Section VI-C.
  • In Section VI, we will classify existing protocols for agent-agent communication. Then, we discuss the risks in Section VI-B and defenses in Section VI-C. Please note that the proposed defenses against compromised agents in this section are also applicable to the risks in Section V-C and Section VII-C.
  • In Section VII, we first show the related protocols in Section VII-A. Then, the risks in this process are also classified into two categories: risks from malicious environments against benign agents (Section VII-B) and risks from compromised agents against benign environments (Section VII-C). The defenses for compromised agents can be referred to Section VI-C. As a result, we only discuss the defenses for the malicious environments in Section VII-D.
Please note that the risks/defenses we focus on are specific for agent communication instead of LLMs. This is because agents and agent communication still have significant differences from LLMs (as we have discussed in Section III). For example, in the agent collaboration process, the accumulation of tiny benign deviations on each step may lead to an intolerable risk in the final result. Miehling et al. [204] also point out that over-focusing on the capabilities of a single model can result in neglecting the emergent behaviors at the system level and underestimating the true risks. As a result, for LLM-targeted attacks that have valuable inspiration for agent security, we will discuss the typical representatives of them instead of listing all papers exhaustively.